OpenStudy (anonymous):

24q7 – 42q4r + 36q3r2 – 63r3 Factor completely

OpenStudy (callisto):

\[24q^7 – 42q^4r + 36q^3r^2 – 63r^3\]\[= 6q^4 (4q^3-7r) + 9r^2(4q^3-7r) \]\[= (6q^4 +9r^2)(4q^3-7r) \]\[= 3(2q^4 +3r^2)(4q^3-7r) \]